3.Shelhorse Corporation produces and sells a single product. Data concerning that product appear below:

Per Unit

Percent of Sales

Selling price

$

210

100

%

Variable expenses

84

40

%

Contribution margin

$

126

60

%

Fixed expenses are $349,000 per month. The company is currently selling 4,400 units per month.

Required: ( 5 points)

The marketing manager believes that a $18,000 increase in the monthly advertising budget would result in a 180 unit increase in monthly sales. What should be the overall effect ($ amount) on the company's monthly net operating income of this change?
Answer #1
Current Proposed
Unit Sales 4400 (4400+180)=4580
Sales (4400*210)=924,000 (4580*210)=961800
Variable expenses (4400*84)=369600 (4580*84)=384720
Contribution margin 554400 577080
Fixed expenses 349,000 (349,000+18000)=367,000
Net operating income 205400 210080

Hence increase in Net operating income=(210080-205400)=$4680.